Problem 11-24A Analyzing journal entries for treasury stock transactions LO 11-4, 11-5

The following correctly prepared entries without explanations pertain to Corners Corporation:

Account Title Debit Credit
1. Cash 1,800,000
Common Stock 1,000,000
Paid-in Capital in Excess of Par Value 800,000
2. Treasury Stock 44,800
Cash 44,800
3. Cash 38,000
Treasury Stock 35,200
Paid-in Capital in Excess of Cost of Treasury Stock 2,800

The original sale (Entry 1) was for 400,000 shares, and the treasury stock was acquired for $8 per share (Entry 2).

PLEASE SHOW ALL WORK

Problem 11-24A Analyzing journal entries for treasury stock transactions LO 11-4,11-5 The following correctly prepared entries without explanations pertain to Corners Corporation Account Title Debit 1,800,000 Credit 1. Cash Common Stock Paid-in Capital in Excess of Par Value 1,000,000 800,000 2. Treasury Stock Cash 44,800 44,800 3. Cash 38,000 Treasury Stochk Paid-in Capital in Excess of Cost of Treasury Stock 35,200 2,800 The original sale (Entry 1) was for 400,000 shares, and the treasury stock was acquired for $8 per share (Entry 2) Required a. What was the sales price per share of the original stock issue? (Round your answer to 2 decimal places.) Sales price per share b. How many shares of stock did the corporation acquire in Entry 2? Corporation acquired shares of stock
c. How many shares were reissued in Entry 3? shares hares reissued d. How many shares are outstanding immediately following Entries 2 and 3, respectively? Outstanding Shares Entry 2 Entry 3
Answer #1

a)

400,000 shares were issued for a total sum of $1,800,000

Hence, issue price of 1 share = 1,800,000/400,000

= $4.5

b)

Treasury shares were acquired for a total sum of $44,800 at $8 per share.

Hence, number of shares acquired = 44,800/8

= 5,600

c)

When treasury shares are re-issued, treasury stock is credited by the cost price paid to acquire the treasury shares.

Hence, number of shares re-issued in entry 3 = 35,200/8

= 4,400

d)

Outstanding shares after entry 2 = Number of shares issued - Number of shares acquired

= 400,000 - 5,600

= 394,400

Outstanding shares after entry 3 = Number of shares issued - Number of shares acquired + Number of shares re-issued

= 400,000 - 5,600 + 4,400

= 398,800

Kindly comment if you need further assistance. Thanks